Barney decides to quit his job as a corporate accountant, which pays $14,000 a month, and goes into business for himself as a certified public accountant.

He runs his business from his converted garage apartment, which he could rent out for $320 a month if he wasn’t using it as a home office. He must purchase office supplies worth $70 a month, and his monthly electricity bill has increased by $50 now that he is working out of his home office.

After six months of working from home, Barney has earned an average of $16,000 per month.

Instructions: Enter your answers as a whole number.

a. What are Barney’s average monthly accounting profits?

     $   

b. What are Barney’s average monthly economic profits?

     $   
Answer #1

a) Average monthly accounting profits

= TR-TC

= (16000-120) = 15880

b) Average monthly economic profits

= TR-TC-(Opportunity costs)

= 15880-(14000+320)

= 1560